OpenStudy (anonymous):

simplify \[4\sqrt[4]{5} + 7\sqrt[3]{5} - 9\sqrt[3]{5} + 6\sqrt[4]{5}\] simplify \[4\sqrt[4]{5} + 7\sqrt[3]{5} - 9\sqrt[3]{5} + 6\sqrt[4]{5}\] @Mathematics

OpenStudy (anonymous):

\[10\sqrt[4]{5}-2\sqrt[3]{5}\]

myininaya (myininaya):

just pretend you have 4x+7y-9y+6x

myininaya (myininaya):

you have like terms right?

OpenStudy (anonymous):

Then collect like terms

myininaya (myininaya):

4x+6x+7y-9y

myininaya (myininaya):

10x-2y

myininaya (myininaya):

see our x is moneybird's \[\sqrt[4]{5}\] and our y is moneybird;s \[\sqrt[3]{5}\]